# Transformer Copilot: Learning from The Mistake Log in LLM Fine-tuning

### v1

@@ 7 Conclusion
Appendix A Broader impact and Limitation
Broader Impact.
Limitation.
Appendix B Additional Details on Transformer Copilot
B.1 Architectural Advantages of T-Copilot.
@@ Appendix B Additional Details on Transformer Copilot
B.2 Decoder-only Copilot Details
@@
Appendix C Proof of Theorem 4.1
@@
$\epsilon_{P}^{2}$ $:=\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}\left[\left(p_{t,i}[k]-\mathbb{E}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}[k]\mid\hat{y}_{t,<i}]\right)^{2}\right]<\infty,$
$\sigma_{P}^{2}$ $:=\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}\big{[}\text{Var}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}[k]\mid\hat{y}_{t,<i}]\big{]}<\infty,$
$\epsilon_{C}^{2}$ $:=\mathbb{E}_{\begin{subarray}{c}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}\big{[}\big{(}p_{t,i}[k]-\hat{p}_{t,i}[k]-\mathbb{E}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}[k]\mid f^{C}_{t,<i}]\big{)}^{2}\mid\hat{y}_{t,<i}\big{]}<\infty,$
$\sigma_{C}^{2}$ $:=\mathbb{E}_{\begin{subarray}{c}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}\big{[}\text{Var}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}[k]\mid f^{C}_{t,<i}]\mid\hat{y}_{t,<i}\big{]}<\infty$
@@
Firstly, by the law of total expectation w.r.t.$(X_{t},Y_{t})$ and the bias-variance decomposition w.r.t. $\hat{p}_{t,i}$,
@@
[…]
@@
[…]
@@
[…]
Together, it follows that
@@
[…]
@@
[…]
@@
[…]
@@
[…]
which is strictly smaller than 00 as long as
$0<\lambda<\frac{2\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}\big{(}\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}-\epsilon_{C}\big{)}}{\big{(}\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}-\epsilon_{C}\big{)}^{2}+\sigma_{C}^{2}}=:\lambda_{0}.$
Appendix D Additional Empirical Analysis
@@
In our empirical analysis, we choose LLaMA-3.2-3B as the Pilot model and T-Copilot-1B as the Copilot model. The Copilot model’s implementation details are the same as stated in Appendix F.2. We evaluate on two reasoning tasks, including SIQA [71] and MAWPS [43]. The dataset details are provided later in Appendix E.
@@
Appendix E Datasets
E.1 Commonsense Reasoning
@@ E.2 Arithmetic Reasoning
E.3 Downstream tasks: Recommendation
@@
E.4 Fine-tuning Dataset Template
@@
Appendix F Experiment Setups
F.1 Hyperparameters and Training/Inference Details
Tables 4-9 present our hyperparameter settings of each task for reproducibility. We perform hyperparameter tuning for both T-Copilot and baseline methods. Unless otherwise specified, both our method and baseline implementations use beam search decoding [23] during inference. All experiments have been run three times with random seeds, reporting average accuracy. For FLAN-T5, LLaMA-3, and Qwen2.5 models, checkpoints are saved every 1,000 steps to track parameters and monitor training to ensure robustness and avoid overfitting.
@@ F.1 Hyperparameters and Training/Inference Details
F.2 T-Copilot Configurations and Implementations
@@ F.3 Baseline Details
[…]
@@
Appendix G Additional Experiments
G.1 Full Table Report on Baseline Comparison
Table 11: Full performance comparison (%) with frontier baselines under matched-parameter scales. Results are averaged over 3 independent runs.
@@ G.1 Full Table Report on Baseline Comparison
$>$8B-level Frontier LLMs
@@
G.2 Downstream Recommendation Evaluation
@@
[…]
G.3 Transferability
T-Copilot-1B HellaSwag BoolQ GSM8K SVAMP Overall Impr.
with new LLaMA-3.2-1B 63.1 65.2 32.2 51.4 $\uparrow$ 6.1%
with new LLaMA-3.3-3B 91.4 70.2 58.8 68.5 $\uparrow$ 4.2%
with new LLaMA-3.1-8B 93.1 71.7 66.0 75.8 $\uparrow$ 2.4%
In the T-Copilot learning framework, the Copilot model is fine-tuned alongside but separately from the Pilot model. Since the same type of models generally have similar learning trajectories under identical training settings, we further investigate: Can the Copilot model leverage the mistake log of one Pilot model and still be effective on another Pilot model of the same type?
We conduct controlled experiments on LLaMA-3 series models in which we directly apply a finetuned 1B Copilot model to new Pilot models during inference. The new Pilot model shares the same architecture as the original one but is trained independently. Note that the Copilot model does not "see" or "learn" any information from the new Pilot model, as they are not jointly trained during finetuning. In Table 14, transferring the Copilot model leads to a slight $\pm$0.2performance difference compared to applying the Copilot to the initial Pilot models (jointly training together). We hypothesize that the minor discrepancy is due to the hardware inference differences between the original and new Pilot models. Nonetheless, the transferred Copilot model still delivers substantial performance gains for the new Pilot and consistently outperforms competing baselines. These results demonstrate that T-Copilot’s error-correction capabilities are not tightly coupled to a specific Pilot model and can be effectively transferred without additional rounds of fine-tuning.
G.4 Scalability
Figure 7 illustrates the relationship between accuracy and model parameter size for T-Copilot. Overall, incorporating the Copilot model consistently demonstrates improved performance as model size increases. We analyze the relationship between performance accuracy ($A$) and model parameter size ($N$) in billions. The derived equations for our method are as follows: for Flan-T5 backbones: $A\simeq 8.74\cdot\log_{10}(N)+40.17$; For LLaMA-3 backbones: $A\simeq 29.58\cdot\log_{10}(N)+50.20$, and for Qwen2.5 backbones: $A\simeq 12.40\cdot\log_{10}(N)+71.80$.
G.5 Efficiency Analysis on Transformer Copilot
@@
Buffer of the Mistake Log. As described in Section 3, we maintain a Mistake Log to record the Pilot model’s internal learning signals, which serve as training data for the Copilot model. To store this information efficiently with minimal GPU and CPU memory overhead, we detach all relevant outputs from the Pilot’s forward pass and store them in a CPU-resident buffer. By default, we use a fixed-size buffer that retains the most recent 128 training rounds. The buffer is updated at each training step, and all Copilot training samples are drawn exclusively from it. This design keeps the additional memory footprint lightweight, typically under 500MB of the CPU memory and less than 200MB of the GPU memory.
G.6 Ablations on Transformer Copilot
@@
Appendix H Additional Related Works
@@
Transformers for Language Modeling The Transformer is a sequence-to-sequence model architecture that employs attention-based mechanisms, making it highly effective for autoregressive language modeling [2, 13, 65]. The vanilla Transformer [78] follows an encoder-decoder structure, comprising a stack of identical layers in both the encoder and decoder components. Each layer consists of a multi-head self-attention mechanism [15], layer normalization [4], a position-wise feedforward network, and residual connection [29]. The encoder-decoder structure serves as the foundation for many early-stage influential LLMs, such as T5 [66], and BART [45]. These models have demonstrated strong capabilities on certain generation tasks [39, 67].
On the other hand, (causal) decoder-only Transformer models [63, 64], trained with the autoregressive language modeling objective [70, 79], have demonstrated exceptional performance in open-ended generation and reasoning tasks [13, 83, 82]. The superior generalization capabilities have established decoder-only Transformers as the backbone of recent state-of-the-art LLMs such as PaLM [16], Falcon [3], LLaMA [77, 21], and ChatGPT [1, 51, 36].
In this work, we develop the Transformer Copilot framework to support both encoder-decoder and decoder-only Transformer architectures. Our intuition is to provide flexibility across a broad range of model configurations and downstream task scenarios.
@@
LLMs Adaptation with Fine-tuning. Large language models perform well across many NLP tasks, yet adapting them to specialized tasks remains challenging [81]. The standard solution, full-parameter fine-tuning [52], retrains all model parameters on task-specific data. Applying full fine-tuning has proven effective at improving performance, but can sometimes be computationally costly [74]. Recent work on parameter-efficient fine-tuning approaches [40, 32, 86, 98, 28], such as prefix-tuning [47] and LoRA [32], aims to reduce the computational overhead by tuning only a small subset of model parameters while still leveraging the expressive power of pre-trained models. Our learning framework builds upon the aforementioned methods’ fine-tuning paradigm and aims to refine the fine-tuning and inference by utilizing mistake information during the models’ learning trajectory. Additionally, since the Copilot model retains the decoder module structure, our framework can seamlessly integrate with various adaptation techniques such as DoRA [50] and ReFT [85].

### v2

@@ Acknowledgment
Appendix A Additional Details on Transformer Copilot
A.1 Architectural Advantages of T-Copilot.
@@ Appendix A Additional Details on Transformer Copilot
A.2 Decoder-only Copilot Details
@@
A.3 Buffer of the Mistake Log
As described in Section 3, we maintain a Mistake Log to record the Pilot model’s internal learning signals, which serve as training data for the Copilot model. To store this information efficiently with minimal GPU and CPU memory overhead, we detach all relevant outputs from the Pilot’s forward pass and store them in a CPU-resident buffer. By default, we use a fixed-size buffer that retains the most recent 128 training rounds. The buffer is updated at each training step, and all Copilot training samples are drawn exclusively from it. This design keeps the additional memory footprint lightweight, typically under 500MB of the CPU memory and less than 200MB of the GPU memory.
A.4 Conceptual Rationale on T-Copilot’s Learning Objective
We elaborate here on the motivation for adopting the RMSE loss in training T-Copilot. Recall that the objective of the Copilot is to predict the distribution discrepancy, defined as
$\ell_{t}(p_{t,i},\hat{p}_{t,i})=p_{t,i}-\hat{p}_{t,i},$
in our formulation. Note that $\ell_{t}(p_{t,i},\hat{p}_{t,i})$ is not a valid probability distribution, since $\sum_{i}\ell_{t}(p_{t,i},\hat{p}_{t,i})\neq 1$. Therefore, it is natural to formulate this task as a regression problem, for which the RMSE loss is commonly adopted.
In contrast, common language modeling objectives such as Cross-Entropy (CE) and KL Divergence are designed for distribution fitting and are not directly applicable in this setting, unless we manipulate $\ell_{t}(p_{t,i},\hat{p}_{t,i})$ with a softmax function to resemble a valid distribution. In that case, however, the loss would no longer optimize the original discrepancy (i.e., $p_{t,i}-\hat{p}_{t,i}$), leading to potential information loss. In Appendix F.4, we further emprically validate the effectiveness of the RMSE loss for T-Copilot training.
Appendix B Proof of Theorem 4.1
@@ A.4 Conceptual Rationale on T-Copilot’s Learning Objective
$\epsilon_{P}^{2}$ $=\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}\left[\left(p_{t,i}[k]-\mathbb{E}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}[k]\mid\hat{y}_{t,<i}]\right)^{2}\right]<\infty,$
$\sigma_{P}^{2}$ $=\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}\big[\text{Var}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}[k]\mid\hat{y}_{t,<i}]\big]<\infty,$
$\epsilon_{C}^{2}$ $=\mathbb{E}_{\begin{subarray}{c}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}\big[\big(p_{t,i}[k]-\hat{p}_{t,i}[k]-\mathbb{E}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}[k]\mid f^{C}_{t,<i}]\big)^{2}\mid\hat{y}_{t,<i}\big]<\infty,$
$\sigma_{C}^{2}$ $=\mathbb{E}_{\begin{subarray}{c}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}\big[\text{Var}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}[k]\mid f^{C}_{t,<i}]\mid\hat{y}_{t,<i}\big]<\infty$
@@
[…]
@@
$={}$ $\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}[\mathbb{E}_{\theta^{P}\sim A^{P}}[(p_{t,i}-\hat{p}_{t,i})^{2}]]$
$={}$ $\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}[(p_{t,i}-\mathbb{E}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}])^{2}+\operatorname{Var}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}]]$
$={}$ $\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}[(p_{t,i}-\mathbb{E}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}])^{2}]+\mathbb{E}_{(X_{t},Y_{t})\sim\mathcal{D}}[\operatorname{Var}_{\theta^{P}\sim A^{P}}[\hat{p}_{t,i}]]$
$={}$ $\epsilon_{P}^{2}+\sigma_{P}^{2}.$
@@
$={}$ $\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[\mathbb{E}_{\theta^{C}\sim A^{C}}[(p_{t,i}-\hat{p}_{t,i}-f^{C}_{t,i})^{2}]]$
$={}$ $\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i}-\mathbb{E}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}])^{2}+\operatorname{Var}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}]]$
$={}$ $\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i}-\mathbb{E}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}])^{2}]+\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[\operatorname{Var}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}]]$
$={}$ $\epsilon_{C}^{2}+\sigma_{C}^{2}.$
@@
$={}$ $\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[\mathbb{E}_{\theta^{C}\sim A^{C}}[(p_{t,i}-\hat{p}_{t,i})(p_{t,i}-\hat{p}_{t,i}-f^{C}_{t,i})]]$
$={}$ $\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i})(p_{t,i}-\hat{p}_{t,i}-\mathbb{E}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}])]$
$\leq{}$ $\sqrt{\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i})^{2}]\cdot\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i}-\mathbb{E}_{\theta^{C}\sim A^{C}}[f^{C}_{t,i}])^{2}]}$
$={}$ $\sqrt{(\epsilon_{P}^{2}+\sigma_{P}^{2})\cdot\epsilon_{C}^{2}}=\epsilon_{C}\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}.$
Together, as long as $0\leq\lambda\leq 1$,
@@
$={}$ $\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ \theta^{C}\sim A^{C}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[((1-\lambda)(p_{t,i}-\hat{p}_{t,i})+\lambda(p_{t,i}-\hat{p}_{t,i}-f^{C}_{t,i}))^{2}]-\mathbb{E}[(p_{t,i}-\hat{p}_{t,i})^{2}]$
$={}$ $(1-\lambda)^{2}\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ \theta^{C}\sim A^{C}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i})^{2}]+\lambda^{2}\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ \theta^{C}\sim A^{C}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i}-f^{C}_{t,i})^{2}]$
@@
$={}$ $(1-\lambda)^{2}\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i})^{2}]+\lambda^{2}\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ \theta^{C}\sim A^{C}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i}-f^{C}_{t,i})^{2}]$
@@
$={}$ $((1-\lambda)^{2}-1)\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i})^{2}]+\lambda^{2}\mathbb{E}_{\begin{subarray}{l}\theta^{P}\sim A^{P}\\ \theta^{C}\sim A^{C}\\ (X_{t},Y_{t})\sim\mathcal{D}\end{subarray}}[(p_{t,i}-\hat{p}_{t,i}-f^{C}_{t,i})^{2}]$
@@
$\leq{}$ $((1-\lambda)^{2}-1)(\epsilon_{P}^{2}+\sigma_{P}^{2})+\lambda^{2}(\epsilon_{C}^{2}+\sigma_{C}^{2})+2(1-\lambda)\lambda\epsilon_{C}\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}$
$={}$ $\lambda\big(\big(\big(\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}-\epsilon_{C}\big)^{2}+\sigma_{C}^{2}\big)\lambda-2\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}\big(\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}-\epsilon_{C}\big)\big),$
which is strictly smaller than 0 as long as
$0<\lambda<\min\bigg\{1,\frac{2\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}\big(\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}-\epsilon_{C}\big)}{\big(\sqrt{\epsilon_{P}^{2}+\sigma_{P}^{2}}-\epsilon_{C}\big)^{2}+\sigma_{C}^{2}}\bigg\}=:\lambda_{0}.$
Appendix C Additional Empirical Analysis
@@
In our empirical analysis, we choose LLaMA-3.2-3B as the Pilot model and T-Copilot-1B as the Copilot model. The Copilot model’s implementation details are the same as stated in Appendix E.2. We evaluate on two reasoning tasks, including SIQA [70] and MAWPS [42]. The dataset details are provided later in Appendix D.
@@
Appendix D Datasets
D.1 Commonsense Reasoning
@@ D.2 Arithmetic Reasoning
D.3 Downstream tasks: Recommendation
@@
D.4 Fine-tuning Dataset Template
@@
Appendix E Experiment Setups
E.1 Hyperparameters and Training/Inference Details
Tables 5-10 present our hyperparameter settings of each task for reproducibility. We perform hyperparameter tuning for both T-Copilot and baseline methods. Unless otherwise specified, both our method and baseline implementations use beam search decoding [23] during inference. All experiments have been run three times with random seeds, reporting average accuracy. For FLAN-T5, LLaMA-3, and Qwen2.5 models, checkpoints are saved every 1,000 steps to track parameters and monitor training to ensure robustness and avoid overfitting.
@@ E.1 Hyperparameters and Training/Inference Details
E.2 T-Copilot Configurations and Implementations
@@ E.3 Baseline Details
(iii) TIES [89]: $\text{T5-small}_{12}$ and $\text{T5-base}_{24}$ are T5 type models merged using the TIES method. $\text{T5-small}_{12}$ merges two T5-small models and extends the original T5-small to 12 encoder and decoder layers. And $\text{T5-base}_{24}$ merges two T5-base models and extends the original T5-base to 24 encoder and decoder layers by duplicating existing layers.
@@
Appendix F Additional Experiments
F.1 Full Table Report on Baseline Comparison
Table 12: Full performance comparison (%) with frontier baselines under matched‐parameter scales. Results are averaged over 3 independent runs.
@@ F.1 Full Table Report on Baseline Comparison
>>8B-level Frontier LLMs
@@
F.2 Downstream Recommendation Evaluation
@@
In Table 13 and Table 14, we report the results of T-Copilot on two downstream recommendation datasets: Beauty and LastFM. We choose T5 and LLaMA-3 series models as the backbone Pilot models. Overall, T-Copilot improves the Pilot models by an average of 16.6% across all evaluation metrics on the two datasets. Furthermore, compared to other baselines, incorporating T-Copilot achieves 16.7% and 8.6% higher performance than $\text{T5-small}_{12}$ and $\text{T5-base}_{24}$, respectively, on Beauty and LastFM, while using 30M and 126M fewer parameters. These results demonstrate that the error-correction capabilities of T-Copilot are not confined to reasoning tasks but also generalize effectively to other application domains, such as recommendation, where precise LLM decision-making is critical for downstream utility.
F.3 Efficiency Analysis on Transformer Copilot
@@
F.4 Ablations and Analyses on Transformer Copilot
@@
Table 19: Comparison of pooling methods for LLaMA-3.1-8B + T-Copilot-1B.
Pooling Methods GSM8K AQuA
Max Pooling 63.2 37.5
Sum Pooling 65.8 38.3
Mean Pooling 66.1 38.9
Table 20: Comparison of loss functions for Qwen2.5-7B + T-Copilot-3B.
Loss Type PIQA AQuA
CE 88.1 60.7
KL Divergence 90.4 62.8
RMSE 92.5 64.2
Pooling Methods of Copilot’s Attention Input. In Section 3, we refine the Copilot model’s attention mechanism by incorporating the mean-pooled hidden representations from the Pilot model as inputs (Eq. 6). To assess the effect of this design choice, we also experiment with alternative pooling strategies, including max pooling and sum pooling. The comparative results are presented in Table 20. From the results, the mean pooling empirically provides the best balance of stability and efficiency.
Loss Types of T-Copilot. We adopt the RMSE loss as the training objective for T-Copilot. To validate this choice, we conduct ablations with alternative loss formulations, including Cross-Entropy and KL Divergence. The results, summarized in Table 20, show that RMSE consistently yields superior performance across benchmarks.
Appendix G Additional Related Works
@@
Transformers for Language Modeling The Transformer is a sequence-to-sequence model architecture that employs attention-based mechanisms, making it highly effective for autoregressive language modeling [2, 13, 64]. The vanilla Transformer [77] follows an encoder-decoder structure, comprising a stack of identical layers in both the encoder and decoder components. Each layer consists of a multi-head self-attention mechanism [15], layer normalization [4], a position-wise feedforward network, and residual connection [29]. The encoder-decoder structure serves as the foundation for many early-stage influential LLMs, such as T5 [65], and BART [44]. These models have demonstrated strong capabilities on certain generation tasks [38, 66]. On the other hand, (causal) decoder-only Transformer models [62, 63], trained with the autoregressive language modeling objective [69, 78], have demonstrated exceptional performance in open-ended generation and reasoning tasks [13, 82, 81]. The superior generalization capabilities have established decoder-only Transformers as the backbone of recent state-of-the-art LLMs such as PaLM [16], Falcon [3], LLaMA [76, 21], and ChatGPT [1, 50, 35]. In this work, we develop the Transformer Copilot framework to support both encoder-decoder and decoder-only Transformer architectures. Our intuition is to provide flexibility across a broad range of model configurations and downstream task scenarios.
@@
LLMs Adaptation with Fine-tuning. Large language models perform well across many NLP tasks [98, 92], yet adapting them to specialized tasks remains challenging [80]. The standard solution, full-parameter fine-tuning [51], retrains all model parameters on task-specific data. Applying full fine-tuning has proven effective at improving performance, but can sometimes be computationally costly [73]. Recent work on parameter-efficient fine-tuning approaches [39, 32, 86, 100, 28], such as prefix-tuning [46] and LoRA [32], aims to reduce the computational overhead by tuning only a small subset of model parameters while still leveraging the expressive power of pre-trained models. Our learning framework builds upon the aforementioned methods’ fine-tuning paradigm and aims to refine the fine-tuning and inference by utilizing mistake information during the models’ learning trajectory. Additionally, since the Copilot model retains the decoder module structure, our framework can seamlessly integrate with various adaptation techniques such as DoRA [49] and ReFT [85].